OpenStudy (anonymous):

check my complex arithmetic? evaluate: final result in polar form "*" means conjugate ((-.5 + i(sqrt(3)/2))*)^8 = =(-.5 - i(sqrt(3)/2)^8 =(e^(i*2pi/3))^8 =e^(8 + i*2pi/3) is this correct?

OpenStudy (anonymous):

\[z^8=\left(-\frac{1}{2}+i\frac{\sqrt3}{2}\right)^8~~?\]

OpenStudy (anonymous):

I just found the equation button 1 sec

OpenStudy (anonymous):

I just have \[z = ((-\frac{ 1 }{ 2 } + i \frac{ \sqrt3 }{ 2 })^*)^8\] where * means the conjugate

OpenStudy (anonymous):

then I get \[= (-\frac{ 1 }{ 2 } - i \frac{ \sqrt3 }{ 2 })^8\] \[=(e ^{i \frac{ 2\pi }{ 3 }})^8\] \[=e ^{8+i \frac{ 2\pi }{ 3 }}\]

OpenStudy (anonymous):

Oh I see. \[\begin{align*}\left(\left(-\frac{1}{2}+i\frac{\sqrt3}{2}\right)^*\right)^8&=\left(-\frac{1}{2}-i\frac{\sqrt3}{2}\right)^8\\\\ &=\left(-\frac{1}{2}-i\frac{\sqrt3}{2}\right)^8\\\\ &=\left(\frac{1}{2}+i\frac{\sqrt3}{2}\right)^8\\\\ &=\left(\cos\frac{\pi}{3}+i\sin\frac{\pi}{3}\right)^8\\\\ &=\cos\frac{8\pi}{3}+i\sin\frac{8\pi}{3}&\text{(by DeMoivre's theorem)}\\\\ &=e^\cdots \end{align*}\]

OpenStudy (anonymous):

so in your 3rd line you can just get rid of the negatives? and they become positive?

OpenStudy (anonymous):

I dont see where they went?

OpenStudy (anonymous):

It was a manner of factoring. For example, \[(-a-b)^2=\bigg((-1)(a+b)\bigg)^2=(-1)^2(a+b)^2=(a+b)^2\] This is true for any even power.

OpenStudy (anonymous):

oh ok great thanks! I understand now. you are very good at explaining and answering :)

OpenStudy (anonymous):

You're welcome! And thank you!